OpenStudy (anonymous):

–2x + 1 ≤ –11 show all work and how to graph

OpenStudy (anonymous):

-2x <= -11 - 1 then???

OpenStudy (anonymous):

solve that eqn ..

OpenStudy (anonymous):

-2x <= - 11 -1 ok.

OpenStudy (anonymous):

the given eqn is -2x + 1 <= 11 add -1 on both the sides so -2x + 1 - 1 <= -11 - 1 -2x <= -12 got it ???

OpenStudy (anonymous):

sry there must be negative sign in firsst eqn -2x + 1 <= -11

OpenStudy (anonymous):

which step is confusing ????

OpenStudy (anonymous):

yaaa

OpenStudy (anonymous):

but you need to simplify it further

OpenStudy (anonymous):

divide both sides by 2

OpenStudy (anonymous):

yaaa so the answer is -x<= -6 which can be wriiten as x >= 6

OpenStudy (anonymous):

thanks for the help i understand it better now

OpenStudy (anonymous):

ok...:)